# Soliton solutions to the Sawada–Kotera equation

### v1

@@ 6 Soliton solutions to the Sawada–Kotera equation
    Using (6.15) on the right-hand side of (6.16), we express each $\gamma_{j}$ as a product of $r_{j}$ and a constant determined by $\{\eta_{j}\}_{j=1}^{\mathbf{N}}.$ By using the right-hand side of (6.16) on the right-hand side of (5.27), we express all entries of the matrix $\mathbf{M}(x)$ appearing on the right-hand side of (5.25) in terms of the elements in $\{\eta_{j},r_{j}\}_{j=1}^{\mathbf{N}}.$ Using the resulting expression for $\mathbf{M}(x)$ on the right-hand side of (5.45), we obtain the solution $Q$ to (2.1) in terms of the elements in $\{\eta_{j}\}_{j=1}^{\mathbf{N}}.$
@@
    Alternatively, we obtain $Q$ by proceeding as follows. Using either of the two roots specified in (6.15), we uniquely determine the quantity $\Delta(x)$ in terms of the parameters in the set $\{k_{j},r_{j}\}_{j=1}^{\mathbf{N}}$ or equivalently in $\{\eta_{j},r_{j}\}_{j=1}^{\mathbf{N}}.$ Using (6.7) in (5.45) we recover the solution $Q$ to (2.1) as
    $Q(x)=6\,\displaystyle\frac{d}{dx}\left[\displaystyle\frac{\Delta^{\prime}(x)}{\Delta(x)}\right],$ (6.17)
@@
    We recall that $\Delta(x)$ is a real-valued polynomial of degree $\mathbf{N}$ in the $\mathbf{N}$ real variables $\chi_{1},\dots,\chi_{\mathbf{N}}.$ The constant term not containing any of those $\mathbf{N}$ variables is equal to $1.$ In each term, a specific $\chi_{j}$ either does not appear or appears only to the first power. From (5.51), (5.52), (6.1), (6.7)–(6.10) it follows that the coefficient of $\chi_{1}$ is equal to $\alpha_{11}/(2\alpha_{0}),$ where we have
@@
    and hence $\chi_{j}$ remains positive for all real values of $x$ and $t.$ By choosing the sign of each $r_{j}$ appropriately so that the right-hand side of (6.19) is positive, we observe from (6.20) and (6.21) that the quantity $y_{j}$ defined in (6.20) remains positive for $1\leq j\leq\mathbf{N}$ for all $x$ and $t.$ We can then write $\Delta(x)$ in terms of $y_{j}$ as
@@ 7 Soliton behaviors
where $y_{1}$ is obtained from (6.20) with $\chi_{1}$ being the the quantity we get from (6.21) when $j=1.$ The coefficient $\alpha_{11}/(2\alpha_{0})$ in (6.20) is evaluated with the help of (5.54), (6.14), and (6.18). By using (5.54) in (6.14), we obtain
@@
where the first line refers to the case $P\equiv 0$ and the second line refers to the case $P=Q_{x}.$ Next, using (5.54) and (7.2) in (6.20), we obtain the quantity $y_{1}$ as
$y_{1}=\begin{cases}-2r_{1}\chi_{1},\qquad P\equiv 0,\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr r_{1}\chi_{1},\qquad P=Q_{x},\end{cases}$ (7.3)
where we recall that $\chi_{1}$ is the exponential function in (6.21) when $j=1.$ We remark that either of the two lines in (7.3) yields two equivalent values for $y_{1}$ by restricting the parameter $r_{1}$ in the first line to negative values and by restricting the parameter $y_{1}$ in the second line to positive values. Using (7.1) and (7.3) in (6.17), we obtain the $1$-soliton solution to (2.1) as
@@
where it is understood that $r_{1}<0$ in the first line and $r_{1}>0$ in the second line. We remark that the substitution $r_{1}\mapsto-r_{1}/2$ in the first line yields the second line in (7.4).
In Figure 7.1 we show the behavior of the $1$-soliton solution corresponding to the input data set with the parameter values given by
@@
where we use the second line of (7.4) corresponding to $P=Q_{x}.$ By using the method of Section 6, we explicitly construct the $1$-soliton solution with the input data set in (7.5). The $1$-soliton solution $Q(x,t)$ in (7.4) is a real-valued solitary wave moving from the left to the right with the constant speed $9\eta_{1}^{4},$ which is equal to $9$ in this particular case. We remark that the $1$-soliton behavior for the Sawada–Kotera equation (2.1) illustrated in Figure 7.1 is similar to the behavior [1, 2, 11, 12, 23] of the $1$-soliton solution to the KdV equation in the sense that the soliton moves from the left to the right without changing its shape.
Figure 7.1: The snapshots for the $1$-soliton solution $Q(x,t)$ to (2.1) at t=0.5,t=-0.5, $t=0,$ and $t=0.5,$ respectively.
@@
where the first line refers to the case $P\equiv 0$ and the second line refers to the case $P=Q_{x}.$ Next, using (5.54) and (7.7) in (6.18), we obtain the quantity $y_{1}$ appearing in (6.20) and (7.6) as
@@
with $\chi_{1}$ being the exponential function obtained from (6.21) when $j=1.$ As a result of the ordering $0<\eta_{1}<\eta_{2},$ in the first line of (7.8) we use $r_{1}<0$ and in the second line we use $r_{1}>0$ so that the quantity $y_{1}$ remains positive for all real values of $x$ and $t.$ The value of $y_{2}$ is obtained with the help of (7.8) by using the replacement $(\eta_{1},\eta_{2},r_{1})\mapsto(\eta_{2},\eta_{1},r_{2}).$ Thus, we get
@@
with $\chi_{2}$ being the exponential function obtained from (6.21) when $j=2.$ In order to have $y_{2}$ to remain positive for all real values of $x$ and $t,$ as a result of the ordering $0<\eta_{1}<\eta_{2},$ we see that we must use $r_{2}>0$ in the first line of (7.9) and use $r_{2}<0$ in the second line.
@@
$(\eta_{1},\eta_{2},r_{1},r_{2})=\left(1,\displaystyle\frac{11}{10},1,-1\right),$ (7.10)
where we use the second lines in (7.7)–(7.9) corresponding to the case $P=Q_{x}.$ Using the method of Section 6, we explicitly construct the $2$-soliton solution with the input data set related to (7.10). The left plot in Figure 7.2 shows a snapshot when the two solitons are apart from each other and not yet interacting with each other. The next plot shows a snapshot at the beginning of the nonlinear interactions between the two solitons. The third plot shows a snapshot at the end of the nonlinear interactions. Finally, the right plot shows a snapshot when the two solitons are apart from each other and there are no longer any interactions. Both solitons travel from the left to the right. Initially, before the nonlinear interactions start, the taller soliton behind has a higher speed than the shorter soliton in the front. In this case, the initial speeds are $13.18$ and $9,$ respectively. We remark that the speed value $13.18$ corresponds to $9\eta_{2}^{4}$ and the speed value $9$ corresponds to $9\eta_{1}^{4}$ in this case. During the nonlinear interactions, the energy from the soliton behind is transmitted to the soliton in the front. Once the nonlinear interactions are over, the speed of the taller soliton behind is reduced from $13.18$ to $9$ and its height is reduced from $5.4$ to $4.5.$ On the other hand, the speed of the shorter soliton in the front has increased from $9$ to $13.18,$ and its height has increased from $2.9$ to $3.7.$ The soliton behind is still taller than the soliton in the front, but it is less taller after the nonlinear interactions. The $2$-soliton behavior for the Sawada–Kotera equation observed in this case is different from the $2$-soliton behavior [1, 2, 11, 12, 23] for the KdV equation. For the KdV equation, although both solitons move from the left to the right, each soliton retains its speed and its shape after the nonlinear interactions. For the KdV equation, the taller soliton always has a higher speed than the shorter soliton.
@@
Figure 7.2: The snapshots for the $2$-soliton solution $Q(x,t)$ to (2.1) with the parameter values in (7.10) at t=2,t=-2, t=0.5,t=-0.5, $t=0,$ and $t=2,$ respectively.
In Figure 7.3 we illustrate the behavior of the $2$-soliton solution corresponding to the input data set $\{\eta_{1},\eta_{2},r_{1},r_{2}\}$ with the parameter values given by
$(\eta_{1},\eta_{2},r_{1},r_{2})=\left(\displaystyle\frac{17}{10},\displaystyle\frac{18}{10},1,-1\right),$ (7.11)
where we again use $P=Q_{x}.$ We explicitly construct the $2$-soliton solution with the input data set corresponding to (7.11). As in Figure 7.2, the left plot in Figure 7.3 shows a snapshot when the two solitons are apart from each other and not yet interacting with each other, the next plot shows a snapshot at the beginning of the nonlinear interactions, the third plot shows a snapshot at the end of the nonlinear interactions, and the right plot shows a snapshot when the two solitons are apart from each other and there are no longer any interactions. Again, both solitons travel from the left to the right. Initially, before the nonlinear interactions start, the shorter soliton behind has a higher speed than the taller soliton in the front, with the speeds being $94.48$ and $75.17,$ respectively. We remark that the speed value $94.48$ corresponds to $9\eta_{2}^{4}$ and the speed value $75.17$ corresponds to $9\eta_{1}^{4}.$ During the nonlinear interactions, the energy from the soliton behind is transmitted to the soliton in the front. Once the nonlinear interactions are over, the speed of the shorter soliton behind is reduced from $94.48$ to $75.17$ and its height is reduced from $14.6$ to $13.$ On the other hand, the speed of the taller soliton in the front has increased from $75.17$ to $94.48,$ and its height has increased from $55.1$ to $58.3.$ The soliton behind is still shorter than the soliton in the front. The $2$-soliton behavior for the Sawada–Kotera equation observed in this case is again different from the $2$-soliton behavior for the KdV equation, with the exception of the fact that for both equations all the solitons move from the left to the right.
Figure 7.3: The snapshots for the $2$-soliton solution $Q(x,t)$ to (2.1) with the parameter values in (7.11) at t=0.2,t=-0.2, t=0.08,t=-0.08, $t=0,$ and $t=0.2,$ respectively.
@@
where the coefficients $A_{12},$ $A_{13},$ $A_{23},$ are obtained in terms of $\eta_{1},$ $\eta_{2},$ $\eta_{3}$ by using (6.24). We have
@@
We remark that the values $A_{13}$ and $A_{23}$ can readily be obtained from the value $A_{12}$ by exploiting the symmetrical appearance of $k_{1},$ $k_{2},$ and $k_{3}$ in (5.2). The quantities $y_{1},$ $y_{2},$ $y_{3}$ are as in (6.20) with $\chi_{1},$ $\chi_{2},$ $\chi_{3}$ being the exponential functions appearing in (6.21). Thus, we have
@@
$(\eta_{1},\eta_{2},\eta_{3},r_{1},r_{2},r_{3})=\left(\displaystyle\frac{32}{100},\displaystyle\frac{36}{100},\displaystyle\frac{4}{10},8,-8,8\right),$ (7.13)
where we use the case $P=Q_{x}.$ Via the method of Section 6, we explicitly construct the $3$-soliton solution with the input data set corresponding to (7.13). The left plot in Figure 7.4 shows a snapshot when the three solitons are apart from each other and not yet interacted with each other. The next plot shows a snapshot at the beginning of the nonlinear interactions among the three solitons. The third plot shows a snapshot at the end of the nonlinear interactions. Finally, the right plot shows a snapshot when the three solitons are apart from each other and they no longer interact. All three solitons travel from the left to the right. Initially, before the nonlinear interactions start, the tallest soliton behind has a higher speed than the two shorter solitons in the front. In this case, the initial speeds are $0.23,$ $0.151,$ and $0.094,$ respectively, for the tallest soliton behind, the middle-height soliton in the middle, and the shortest soliton in the front. We remark that the speed value $0.23$ corresponds to $9\eta_{3}^{4},$ the speed value $0.151$ corresponds to $9\eta_{2}^{4},$ and the speed value $0.094$ corresponds to $9\eta_{1}^{4}.$ During the nonlinear interactions, the energy from each soliton behind is transmitted to the next soliton in the front. Once the nonlinear interactions are over, the three solitons are again aligned so that the tallest soliton is still behind, the middle-height soliton is still in the middle, and the shortest soliton is still in the front. However, the speed of the tallest soliton has decreased from $0.23$ to $0.094,$ the speed of the middle-height soliton remains unchanged at the speed value $0.151,$ and the speed of the shortest soliton has increased from $0.094$ to $0.23.$ On the other hand, the height of the tallest soliton behind has decreased from $0.72$ to $0.46,$ the height of the middle-height soliton in the middle has decreased from $0.1$ to $0.065,$ and the height of the shortest soliton in the front has decreased from $0.055$ to $0.017.$ The $3$-soliton behavior for the Sawada–Kotera equation observed in this case is different from the $3$-soliton behavior for the KdV equation. For the KdV equation, although all three solitons move from the left to the right, each soliton retains its speed and its shape after the nonlinear interactions, and a taller soliton always has a higher speed than a shorter soliton.
Figure 7.4: The snapshots for the $3$-soliton solution $Q(x,t)$ to (2.1) with the parameter values in (7.13) at t=300,t=-300, t=100,t=-100, $t=0,$ and $t=300,$ respectively.
$(\eta_{1},\eta_{2},\eta_{3},r_{1},r_{2},r_{3})=\left(1,\displaystyle\frac{11}{10},\displaystyle\frac{12}{10},\displaystyle\frac{1}{10},-\displaystyle\frac{1}{20},1\right),$ (7.14)
@@
where we again use the case $P=Q_{x}.$ We explicitly construct the $3$-soliton solution with the input data set corresponding to (7.14). The left plot in Figure 7.5 shows a snapshot when the three solitons are apart from each other and not yet interacting with each other. The next plot shows a snapshot at the beginning of the nonlinear interactions. The third plot shows a snapshot at the end of the nonlinear interactions. Finally, the right plot shows a snapshot when the three solitons are apart from each other and they no longer interact. All three solitons travel from the left to the right. Initially, before the nonlinear interactions start, the tallest soliton in the front has a lower speed than the middle-height soliton behind and the shortest soliton in the middle. In this case, the initial speeds are $18.66,$ $13.18,$ and $9,$ respectively, for the middle-height soliton behind, the shortest soliton in the middle, and the tallest soliton in the front. During the nonlinear interactions, the energy from each soliton behind is transmitted to the soliton in the front. Once the nonlinear interactions are over, the three solitons are again aligned so that the tallest soliton is still in the front, the shortest soliton is still in the middle, and the middle-height soliton is still behind. However, the speed of the tallest soliton in the front has increased from $9$ to $18.66,$ the speed of the shortest soliton in the middle has remained unchanged at the value $13.18,$ and the speed of the middle-height soliton behind has decreased from $18.66$ to $9.$ On the other hand, the height of the tallest soliton in the front has increased from $12.6$ to $19,$ the height of the shortest soliton in the middle has decreased from $5.3$ to $3.7,$ and the height of the middle-height soliton behind has decreased from $6.5$ to $4.5.$ The behavior of the $3$-soliton solution to the Sawada–Kotera equation observed in this case is different from the $3$-soliton behavior for the KdV equation. For the KdV equation, although all three solitons move from the left to the right, each soliton retains its speed and its shape after the nonlinear interactions, and a taller soliton always has a higher speed than a shorter soliton.
@@
Figure 7.5: The snapshots for the $3$-soliton solution $Q(x,t)$ to (2.1) with the parameter values in (7.14) at t=2,t=-2, t=0.5,t=-0.5, $t=0,$ and $t=2,$ respectively.

### v2

@@ 6 Soliton solutions to the Sawada–Kotera equation
    Using (6.15) on the right-hand side of (6.16), we express each $\gamma_{j}$ as a product of $r_{j}$ and a constant determined by the elements of the set $\{\eta_{j}\}_{j=1}^{\mathbf{N}}.$ By using the right-hand side of (6.16) on the right-hand side of (5.27), we express all entries of the matrix $\mathbf{M}(x)$ appearing on the right-hand side of (5.25) in terms of the elements in $\{\eta_{j},r_{j}\}_{j=1}^{\mathbf{N}}.$ Using the resulting expression for $\mathbf{M}(x)$ on the right-hand side of (5.45), we obtain the solution $Q$ to (2.1) in terms of the elements in $\{\eta_{j}\}_{j=1}^{\mathbf{N}}.$
@@
    Alternatively, we obtain $Q$ by proceeding as follows. Using either of the two roots specified in (6.15), we uniquely determine the quantity $\Delta(x)$ in terms of the parameters in the set $\{k_{j},r_{j}\}_{j=1}^{\mathbf{N}}$ or equivalently in the set $\{\eta_{j},r_{j}\}_{j=1}^{\mathbf{N}}.$ Using (6.7) in (5.45) we recover the solution $Q$ to (2.1) as
    $Q(x)=6\,\displaystyle\frac{d}{dx}\left(\displaystyle\frac{\Delta^{\prime}(x)}{\Delta(x)}\right),$ (6.17)
@@
    We recall that $\Delta(x)$ is a real-valued polynomial of degree $\mathbf{N}$ in the $\mathbf{N}$ real variables $\chi_{1},\dots,\chi_{\mathbf{N}}.$ The constant term not containing any of those $\mathbf{N}$ variables is equal to $1.$ In each term of $\Delta(x),$ a specific $\chi_{j}$ either does not appear or appears only to the first power. From (5.51), (5.52), (6.1), (6.7)–(6.10) it follows that the coefficient of $\chi_{1}$ is equal to $\alpha_{11}/(2\alpha_{0}),$ where we have
@@
    and hence $\chi_{j}$ remains positive for all real values of $x$ and $t.$ By choosing the sign of each $r_{j}$ appropriately so that the right-hand side of (6.19) is positive, we observe from (6.20) and (6.21) that the quantity $y_{j}$ defined in (6.20) remains positive for $1\leq j\leq\mathbf{N}$ for all $x$ and $t.$ We can then write $\Delta(x)$ as
@@ 7 Soliton behaviors
where $y_{1}$ is obtained from (6.20) with $\chi_{1}$ being the quantity we get from (6.21) when $j=1,$ i.e. we have
$\chi_{1}=e^{\sqrt{3}\,\eta_{1}(x-9\eta_{1}^{4}t)}.$ (7.2)
The coefficient $\alpha_{11}/(2\alpha_{0})$ in (6.20) is evaluated with the help of (5.54), (6.14), and (6.18). By using (5.54) in (6.14), we obtain
@@
where the first line on the right-hand side refers to the case $P\equiv 0$ and the second line refers to the case $P=Q_{x}.$ Next, using (5.54) and (7.3) in (6.20), we obtain the quantity $y_{1}$ as
$y_{1}=\begin{cases}-2r_{1}\chi_{1},\qquad P\equiv 0,\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr r_{1}\chi_{1},\qquad P=Q_{x}.\end{cases}$ (7.4)
We remark that either of the two lines on the right-hand side of (7.4) yields two equivalent values for $y_{1}$ by restricting the parameter $r_{1}$ in the first line to negative values and by restricting the parameter $r_{1}$ in the second line to positive values. Using (7.1) and (7.4) in (6.17), we obtain the $1$-soliton solution to (2.1) as
@@
where it is understood that $r_{1}<0$ in the first line on the right-hand side and $r_{1}>0$ in the second line. We observe that the substitution $r_{1}\mapsto-r_{1}/2$ in the first line yields the second line in (7.5). Since $\chi_{1}$ appearing in (7.2) is a function of $(x-9\eta_{1}^{4}t),$ from (7.5) we see that the $1$-soliton $Q(x,t)$ moves from the left to the right with the speed $9\eta_{1}^{4}.$
In Figure 7.1 we show the four snapshots at t=2,t=-2, t=1,t=-1, $t=0,$ and $t=1,$ respectively, for the $1$-soliton solution to (2.1) constructed with the input parameters given by
@@
where we use the second line on the right-hand side of (7.5) corresponding to $P=Q_{x}.$ From (7.5) we see that the same soliton solution can be constructed by using the first line of the right-hand side of (7.5) with the input parameters $\eta_{1}$ and $r_{1}$ given by
$(\eta_{1},r_{1})=\left(1,-\displaystyle\frac{1}{2}\right),$
in which case the corresponding $P$ has the value $P(x,t)\equiv 0.$ The $1$-soliton solution $Q(x,t)$ in (7.5) is a real-valued solitary wave moving from the left to the right with the constant speed $9\eta_{1}^{4},$ which is equal to $9$ in this particular case. We remark that the $1$-soliton behavior for the Sawada–Kotera equation (2.1) illustrated in Figure 7.1 is similar to the behavior [1, 2, 11, 12, 23] of the $1$-soliton solution to the KdV equation in the sense that the soliton moves with a constant speed from the left to the right without changing its shape.
Figure 7.1: The snapshots for the $1$-soliton solution $Q(x,t)$ to (2.1) with the parameter values in (7.6) at t=2,t=-2, t=1,t=-1, $t=0,$ and $t=1,$ respectively.
@@
where the first lines on the right-hand sides refer to the case $P\equiv 0$ and the second lines refer to the case $P=Q_{x}.$ Next, using (5.54) and (7.8) in (6.18), we obtain the quantity $y_{1}$ appearing in (6.20) and (7.7) as
@@
with $\chi_{1}$ being the exponential function given in (7.2). As a result of the ordering $0<\eta_{1}<\eta_{2},$ in the first line on the right-hand side of (7.9) we use $r_{1}<0$ and in the second line we use $r_{1}>0$ so that the quantity $y_{1}$ remains positive for all real values of $x$ and $t.$ The value of $y_{2}$ is obtained with the help of (7.9) by using the replacement $(\eta_{1},\eta_{2},r_{1})\mapsto(\eta_{2},\eta_{1},r_{2}).$ Thus, we get
@@
with $\chi_{2}$ being the exponential function obtained from (6.21) when $j=2,$ i.e. we have
$\chi_{2}=e^{\sqrt{3}\,\eta_{2}(x-9\eta_{2}^{4}t)}.$ (7.11)
In order to have $y_{2}$ to remain positive for all real values of $x$ and $t,$ as a result of the ordering $0<\eta_{1}<\eta_{2},$ we see that we must use $r_{2}>0$ in the first line on the right-hand side of (7.10) and use $r_{2}<0$ in the second line.
@@
$(\eta_{1},\eta_{2},r_{1},r_{2})=\left(\displaystyle\frac{1}{\sqrt{3}},\displaystyle\frac{2}{\sqrt{3}},1,-1\right),$ (7.12)
where we use the second lines on the right-hand sides in (7.8)–(7.10) corresponding to the case $P=Q_{x}.$ Using the method of Section 6, we explicitly construct the $2$-soliton solution with the input data set related to (7.12). Using the values specified in (7.12) as input to the key quantity $\Delta(x,t)$ appearing in (7.7), we obtain
$\Delta(x,t)=1+\displaystyle\frac{14}{3}\,\chi_{1}+7\,\chi_{2}+\displaystyle\frac{14}{9}\,\chi_{1}\chi_{2},$ (7.13)
where the quantities $\chi_{1}$ and $\chi_{2}$ are constructed with the help of (7.2) and (7.11) as
$\chi_{1}=e^{x-t},\quad\chi_{2}=e^{2(x-16t)}.$
Using (7.13) in (6.17), we obtain the solution $Q(x,t)$ to (2.1) as
$Q(x,t)=\displaystyle\frac{28\,\chi_{1}+168\,\chi_{2}+280\,\chi_{1}\,\chi_{2}+\displaystyle\frac{1568}{9}\,\chi_{1}^{2}\,\chi_{2}+\displaystyle\frac{196}{3}\,\chi_{1}\,\chi_{2}^{2}}{\left(1+\displaystyle\frac{14}{3}\,\chi_{1}+7\,\chi_{2}+\displaystyle\frac{14}{9}\,\chi_{1}\,\chi_{2}\right)^{2}},$ (7.14)
with the understanding that we have $P(x,t)=Q_{x}(x,t).$ By exploiting the symmetrical appearance of $k_{1}$ and $k_{2}$ in (5.2), we determine that we obtain the same $Q(x,t)$ given in (7.14) with $P(x,t)\equiv 0$ if we use not the input values in (7.12) but instead use the input values given by
$(\eta_{1},\eta_{2},r_{1},r_{2})=\left(\displaystyle\frac{1}{\sqrt{3}},\displaystyle\frac{2}{\sqrt{3}},-\displaystyle\frac{2}{3},\displaystyle\frac{3}{2}\right).$
The left plot in Figure 7.2 shows the snapshot at t=0.6t=-0.6 when the two solitons are apart from each other and not yet interacting with each other. The taller soliton has the speed $16$ and is behind the shorter soliton having the speed $1.$ Both solitons move from the left to the right. The next plot shows the snapshot at t=0.2t=-0.2 soon after the beginning of the nonlinear interactions between the two solitons. The third plot shows the snapshot at $t=0.2$ when the nonlinear interactions are about to end. Finally, the right plot shows the snapshot at $t=0.6$ when the two solitons are apart from each other and there are no longer any interactions. By comparing the four snapshots in Figure 7.2, we observe that, as a result of the nonlinear interactions, the shorter soliton has been pushed backward as the taller soliton overtakes that shorter soliton. During the nonlinear interactions the taller soliton overtakes the shorter soliton. We also remark that, even though we do not show the snapshot at $t=0$ in Figure 7.2, the two solitons at $t=0$ in this case overlap in such a way as if there were a single soliton at that moment. The $2$-soliton behavior for the Sawada–Kotera equation is similar to the $2$-soliton behavior for the KdV equation.
@@
Figure 7.2: The snapshots for the $2$-soliton solution $Q(x,t)$ to (2.1) with the parameter values in (7.12) at t=0.6,t=-0.6, t=0.2,t=-0.2, $t=0.2,$ and $t=0.6,$ respectively.
@@
where the coefficients $A_{12},$ $A_{13},$ $A_{23}$ are obtained in terms of $\eta_{1},$ $\eta_{2},$ $\eta_{3}$ by using (6.24). We have
@@
We remark that the quantities $A_{13}$ and $A_{23}$ can readily be obtained from the quantity $A_{12}$ by exploiting the symmetrical appearance of $k_{1},$ $k_{2},$ and $k_{3}$ in (5.2). The quantities $y_{1},$ $y_{2},$ $y_{3}$ are as in (6.20) with $\chi_{1},$ $\chi_{2},$ $\chi_{3}$ being the exponential functions appearing in (6.21). Thus, we have
@@
$(\eta_{1},\eta_{2},\eta_{3},r_{1},r_{2},r_{3})=\left(\displaystyle\frac{1}{\sqrt{3}},\displaystyle\frac{2}{\sqrt{3}},\displaystyle\frac{3}{\sqrt{3}},1,-1,-1\right),$ (7.16)
where we use the case $P=Q_{x}.$ Via the method of Section 6, we explicitly construct the $3$-soliton solution with the input data set corresponding to (7.16). For this, we first use (7.16) in (7.15) and obtain the corresponding $\Delta(x,t)$ as
$\Delta(x,t)=1+13\,\chi_{1}+95\,\chi_{2}+\displaystyle\frac{1235}{7}\,\chi_{3}+\displaystyle\frac{1235}{21}\,\chi_{1}\,\chi_{2}+\displaystyle\frac{1235}{4}\,\chi_{1}\,\chi_{3}+247\,\chi_{2}\,\chi_{3}+\displaystyle\frac{247}{12}\,\chi_{1}\,\chi_{2}\,\chi_{3},$ (7.17)
where the quantities $\chi_{1},$ $\chi_{2},$ $\chi_{3}$ are obtained with the help of (6.21) and (7.16) as
$\chi_{1}=e^{x-t},\quad\chi_{2}=e^{2(x-16t)},\quad\chi_{3}=e^{3(x-81t)}.$
The $3$-soliton solution $Q(x,t)$ to (2.1) is then constructed by using (7.17) in (6.17).
@@
The left plot in Figure 7.3 shows the snapshot at t=0.6t=-0.6 when the three solitons are apart from each other and not yet interacting with each other. The next plot shows the snapshot at t=0.1t=-0.1 near the beginning of the nonlinear interactions among the three solitons. The third plot shows the snapshot at $t=0.13$ near the end of the nonlinear interactions. Finally, the right plot shows the snapshot at $t=0.4$ when the three solitons are apart from each other and they no longer interact. We also observe the following. All three solitons travel from the left to the right. Initially, before the nonlinear interactions start, they are aligned in such a way that the tallest soliton is behind the other two solitons, the shortest soliton is ahead of the other two, and the middle-height soliton is between the other two. Prior to any nonlinear interactions, the tallest soliton has the speed 81, the middle-height soliton has the speed $16,$ and the shortest soliton has the speed $1.$ During the nonlinear interactions, the tallest soliton overtakes the other two solitons and the middle-height soliton overtakes the shortest soliton. We further observe that the shortest soliton has been pushed backward as a result of the nonlinear interactions. In Figure 7.3 we do not show the snapshot at $t=0,$ but we remark that the three solitons at $t=0$ in this case overlap in such a way as if there were a single soliton at that moment. The $3$-soliton behavior for the Sawada–Kotera equation is similar to the $3$-soliton behavior for the KdV equation.
@@
Figure 7.3: The snapshots for the $3$-soliton solution $Q(x,t)$ to (2.1) with the parameter values in (7.16) at t=0.6,t=-0.6, t=0.1,t=-0.1, $t=0.13,$ and $t=0.4,$ respectively.
We recall that the input data set with the parameter values in (7.16) yields the quantity $\Delta(x,t)$ in (7.17) and the corresponding $3$-soliton solution $Q(x,t)$ to (2.1) via (6.17), where the associated quantity $P(x,t)$ is related to $Q(x,t)$ as $P(x,t)=Q_{x}(x,t).$ By exploiting the symmetrical appearance of $k_{1},$ $k_{2},$ and $k_{3}$ in (5.2), we construct the same $\Delta(x,t)$ in (7.17) and hence the same $Q(x,t)$ with the associated quantity $P(x,t)$ satisfying $P(x,t)\equiv 0$ if we use not the input values in (7.16) but instead use the input values given by
$(\eta_{1},\eta_{2},\eta_{3},r_{1},r_{2},r_{3})=\left(\displaystyle\frac{1}{\sqrt{3}},\displaystyle\frac{2}{\sqrt{3}},\displaystyle\frac{3}{\sqrt{3}},-\displaystyle\frac{3}{4},\displaystyle\frac{5}{2},-\displaystyle\frac{25}{4}\right).$